OpenStudy (anonymous):

I GIVE METALS, and i will be your fan Which of the following is an equivalent representation of 2-4 ? A. 1/8 B. 1/27 C. 1/16 D. 1/9

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@millsemily

Vocaloid (vocaloid):

2^4 = ?

OpenStudy (millsemily):

Or did you mean 2/4..I dunno what you mean by that.

OpenStudy (anonymous):

2^4

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

right, so 2^4 = ?

OpenStudy (anonymous):

ya

Vocaloid (vocaloid):

no, I'm asking you to calculate 2^4 = ???

OpenStudy (anonymous):

idk what it is

Vocaloid (vocaloid):

2^4 = 2*2*2*2 = ?

OpenStudy (anonymous):

8

Vocaloid (vocaloid):

try again.... 2*2*2*2 = ?

Vocaloid (vocaloid):

good, so if 2^4 = 16, then 2^-4 = 1/16

OpenStudy (anonymous):

k thanx